# JuDGE: Benchmarking Judgment Document Generation for Chinese Legal System

### v2

@@ Abstract.
conference: 2025; Under Review;
@@ 1. Introduction
In recent years, the rapid development of large language models (LLMs) has enabled numerous applications in the legal domain (Brown et al., 2020; Touvron et al., 2023; Scao et al., 2022; Yang et al., 2024), including contract analysis, case law research, and legal document drafting (Shu et al., 2024; Yue et al., 2023; Wang et al., 2024b; Su et al., 2024e). Among these applications, the automated generation of judgment documents is particularly promising. This task presents significant challenges, including the accurate capture of factual details, the application of laws and statutes, and the execution of legal reasoning. Moreover, it has the potential to enhance judicial efficiency and reduce legal workloads.
Judgment documents are authoritative court records that encapsulate legal reasoning, procedural details, and final rulings (Kumar et al., 2013; Li et al., 2023a). As essential components of judicial proceedings, these documents serve as authoritative records of legal decisions and reflect the complexity of legal reasoning and statutory interpretation (Zander, 2015). In practice, drafting judgment documents is a highly labor-intensive task that consumes substantial time and resources in courts. Specifically, creating a judgment document requires the judge to gather a large amount of legal information, including relevant statutes, past case precedents, and fundamental legal knowledge. Once the relevant legal information has been collected, it must be systematically organized and integrated with professional legal reasoning to generate a well-structured and legally sound judgment document. Given the inherently time-consuming and labor-intensive nature of drafting judgment documents, there is significant potential for using AI technologies to assist legal professionals. For example, advanced IR techniques can systematically collect and organize the relevant legal information, while generative AI can support the drafting of well-structured legal texts. This combined approach could substantially reduce manual effort and streamline the document creation process.
@@ 2.1. Judgment Document Structure and Notation
Let $\mathcal{F}$ denote the space of fact descriptions extracted from legal case documents. Each fact description $f\in\mathcal{F}$ provides a clear and unbiased summary of the events and circumstances underlying a legal dispute. The corresponding judgment document, denoted by $j$, belongs to the space $\mathcal{J}$. A complete judgment document $j\in\mathcal{J}$ typically comprises several interdependent sections, including but not limited to:
@@
  • Judicial Reasoning: This section details the court’s evaluation of both the factual and legal aspects of the case. It systematically applies relevant laws, legal principles, and judicial precedents to the established facts. The analysis discusses the evidence, the offense’s nature, and the liability’s determination. It ensures that the legal reasoning behind the decision is transparent and well-founded.
@@
The interdependence between these sections ensures the judgment document maintains logical consistency and legal coherence. Each section builds upon the previous one, with the Judicial Analysis Process being generated based on the Fact Description while simultaneously providing the legal foundation for the Judgment Result.
@@ 3. Dataset Construction
This section outlines the construction process of the JuDGE dataset, starting with the collection of criminal case documents and the statute corpus. Next, it describes the pre-processing steps undertaken to ensure high data quality, followed by expert annotations to verify the dataset’s accuracy and consistency. The section then presents key dataset statistics, offering insights into its composition, and concludes with a discussion of the ethical considerations, highlighting our commitment to privacy, transparency, and regular updates.
@@ 3.3. Expert Annotation
While the data pre-processing steps provided an initial quality filter for judgment documents, we further enhanced the JuDGE dataset through expert annotation. To achieve this, we recruited law students to evaluate each judgment document instance based on five critical aspects: document formatting, the correctness of judicial reasoning, appropriate legal references, the accuracy of crime classification, and the reasonableness of the sentencing. Each document was annotated by two independent experts. If both annotators agreed that the document met the quality criteria, it was classified as “usable.” If discrepancies arose between annotators, the document was excluded from the dataset. The annotation team consisted of seven members recruited from prominent law schools. Our compensation plan, which was based on working hours, ensured that the annotation process was incentivized fairly, offering an average hourly wage of 45 CNY, which is significantly higher than the minimum wage requirements in Beijing.
To evaluate the reliability of the annotations, we applied Cohen’s Kappa coefficient in a binary classification context. The analysis, performed on 2,574 annotated instances, yielded a Kappa value of 0.8012, indicating very high inter-annotator agreement. This strong agreement reflects the consistency and reliability of the expert annotations. After applying these filtering standards and expert annotations, we obtained the final dataset with 2,505 legal documents covering 182 different statutes and 142 distinct crime types. This selection process ensured the dataset’s diversity in legal references and crime types while also maintaining its high quality for the benchmark.
@@ 3.4. Dataset Statistics
Table 2 summarizes the core numerical attributes of the JuDGE dataset and its supplementary legal corpora. The dataset contains 2,505 fact-judgment pairs, covering a diverse range of criminal cases with 142 unique charges and 182 distinct criminal law provisions. This diversity suggests that the dataset is highly applicable for modeling various legal scenarios.
The average lengths of the different sections further reflect the dataset’s balanced level of detail. Fact descriptions average around 652 Chinese characters, providing a succinct yet informative account of case details. On a per-document basis, the inclusion of roughly 1.26 charges and 4.31 statutory articles emphasizes each case’s multifaceted legal context. Moreover, the dataset is supplemented by two extensive external corpora: one containing 103,251 judgment documents and another comprising 55,348 statutory articles. These resources provide rich legal background information that is particularly valuable for advanced IR technologies like retrieval-augmented generation approaches.
@@ 4. Evaluation Framework
We compare the Judicial Reasoning and Judgment Result sections of the generated text against those of the reference using two semantics-based metrics [^4] The intuitive significance of the numerical values presented in “METH.” and “BERTS.” is further clarified by a series of case studies, which can be accessed on our dataset’s official websitehttps://github.com/oneal2000/JuDGE:
@@
We opt against traditional metrics like BLEU or ROUGE, as these n-gram-based measures can overlook the complex reasoning structures and nuanced legal vocabulary inherent in real judgments.
@@ 5.1. Baselines
To support future research, we systematically evaluate several baseline approaches on our proposed JuDGE benchmark using various LLMs. These baselines include few-shot in-context learning and fine-tuning techniques to adapt general LLMs to the judgment document generation task. Additionally, we introduce an advanced baseline, Multi-source Retrieval-Augmented Generation (MRAG), which integrates knowledge from both statute and judgment document corpora, establishing a strong baseline performance for future reference.
@@ 5.1.3. Retrieval Augmented Generation
Building upon the RAG paradigm, we explore and propose an RAG-based approach as a baseline for our task. In practical legal applications, generating a complete judgment document requires the integration of extensive legal information, including relevant statutory regulations, historical case precedents, and fundamental legal principles. To address this requirement, we propose the Multi-Source Retrieval-Augmented Generation (MRAG) baseline that integrates external knowledge to enhance the generation process. This baseline divides judgment generation into two phases: Information Collection and Document Generation. In the Information Collection phase, two distinct retrievers are employed to retrieve relevant information from different sources. Specifically, the Law Retriever targets relevant statutory regulations, while the Case Retriever focuses on retrieving relevant judgment documents. In the Document Generation phase, the LLM leverages the fact description from the dataset along with the information gathered in the information collection phase to generate the complete judgment document.
@@
Table 4. Experimental results on multiple baselines across six different LLMs. QW-3B-Base refers to the Qwen-2.5-3B model, while QW-3B-Chat refers to the Qwen-2.5-3B-Instruct model (the same naming convention applies to the 7B models). “METH.” denotes METEOR, “BERTS.” for BERTScore, and “Prec.” for precision. The best results for each LLM are highlighted in bold.
@@
Model Method Prison Fine Recall Prec. F1 Recall Prec. F1 METH. BERTS. METH. BERTS.
@@ 6. Experimental Results
(1) Legal LLMs (e.g., LexiLaw-6B and Hanfei-7B) show notably poor performance when generating judgment documents directly or via few-shot in-context learning. Even with in-context learning, the performance remains weak. A closer examination of the generated outputs reveals that these models often reuse or partially copy the exemplars provided in the prompt, neglecting to incorporate the unique facts of each case. We suspect that this phenomenon arises from catastrophic forgetting. Since these models were continually pre-trained and fine-tuned on large-scale legal data (e.g., legal QA), their parameters have diverged significantly from their initial state, resulting in the loss of previously acquired knowledge and instruction-following capabilities. (2) Few-shot in-context learning benefits larger chat models more significantly. For QW-7B-Chat, introducing a few-shot example in the prompt consistently enhances performance across almost all the metrics. In contrast, QW-3B-Chat sees only marginal gains from in-context learning. This difference highlights that larger models with more capacity generally exhibit stronger in-context learning capabilities. (3) Supervised fine-tuning substantially improves performance. Regardless of the models’ initial capability, supervised fine-tuning (SFT) leads to improvements over direct generation and in-context learning. This trend is observed across both general-purpose models (e.g., QW-3B-Chat and QW-7B-Chat) and legal LLMs (e.g., LexiLaw-6B and Hanfei-7B). Notably, the improvement is particularly large for legal LLMs: while their performance is almost unusable before fine-tuning, it becomes acceptable and sometimes competitive after the task-specific SFT. These findings reinforce the importance of supervised adaptation for certain downstream tasks. (4) Multi-source retrieval-augmented generation (RAG) offers competitive advantages. For instance, MRAG consistently achieves the highest recall, precision, and F1 score in convicting accuracy. However, while MRAG achieves high recall in referencing accuracy, its precision is relatively lower. We attribute this to our top-$k$ retrieval strategy (here $k=10$), which occasionally supplies multiple irrelevant statutes. As indicated by Table 2, only 4.31 criminal statutes are truly relevant per case. Since current LLMs are not fully adept at filtering out misleading articles, this results in a precision drop.
@@ 7.1. Legal Information Retrieval
Legal information retrieval (IR) differs from open-domain IR as it requires the retrieval system to incorporate extensive legal knowledge and consider both semantic and legal perspectives when modeling the text. A prominent research direction is Legal Case Retrieval, which focuses on retrieving past legal cases that are relevant to a given query case, aiding legal professionals in case analysis and decision-making (Su et al., 2023b; Ma et al., 2021, 2023). Early work in this area primarily relied on lexical matching methods to assess similarity (Rosa et al., 2021). Recently, with the development of Dense Retrieval (Karpukhin et al., 2020; Zhan et al., 2021; Su et al., 2023c; Fang et al., 2024; Su et al., 2023a), researchers have begun to explore methods to enhance retrievers with deeper legal understanding via large-scale pretraining on unlabeled legal corpora. For instance, Sailer (Li et al., 2023a) trains the dense retriever through an autoencoder-like approach, where the embedding of a case’s Fact section is used to reconstruct other sections of the document, thereby infusing the fact embeddings with richer contextual information. Caseformer (Su et al., 2023b) proposes an unsupervised contrastive learning approach for automatically measuring similarity between legal documents. The method leverages legal case attributes such as charges and cited legal provisions to automatically generate positive and negative cases relative to a query case, thereby training dense retrieval models. This fully automated process enables the generation of substantial training data from unlabeled corpora, enhancing retrieval performance. Another major direction is Statute Retrieval, which aims to locate and retrieve relevant statutory articles for given a query. For instance, the annual COLIEE competitions (Goebel et al., 2023; Kim et al., 2022; Rabelo et al., 2022; Li et al., 2023b, c) focus on Japanese legal bar exam questions, where the goal is to retrieve relevant statutes from the Japanese Civil Code. Similarly, the AILA (Bhattacharya et al., 2019) targets the Indian legal system, using queries from Supreme Court of India judgments to retrieve relevant statutes. Beyond professional use cases, statute retrieval also caters to non-professional users who may lack formal legal training; for example, STARD (Su et al., 2024a) is specifically designed to address layperson queries, providing plain-language questions paired with relevant statutory content. Beyond retrieval, Legal QA tasks aim to satisfy the public’s need for accessible legal information (Do et al., 2017), such as the French long-form QA dataset LLeQA (Louis et al., 2024) and GerLayQA (Büttner and Habernal, 2024), which contains laymen’s German legal questions paired with authoritative law book paragraphs.
@@ 8. Conclusion
In this paper, we introduced JuDGE, a comprehensive benchmark for evaluating the generation of judgment documents within the legal domain. By formalizing the task and providing the benchmark dataset, we establish a solid foundation for further research in this important yet underexplored area. Through collaboration with legal professionals, we also developed an automated evaluation framework that measures performance across four key dimensions: penalty accuracy, convicting accuracy, referencing accuracy, and similarity to ground truth. Our experimental results reveal that current approaches struggle to produce high-quality judgment documents. Although supervised fine-tuning and our proposed Multi-source RAG approach show improvements, the performance gap indicates that judgment document generation remains challenging and requires further research and innovation. We hope JuDGE will inspire more advanced techniques in judgment document generation, ultimately helping the legal community improve efficiency and reduce manual workloads.

### v3

@@ Abstract.
copyright: acmcopyrightjournalyear: 2025copyright: ccconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3730295isbn: 979-8-4007-1592-1/2025/07ccs: Information systems Retrieval tasks and goals
@@ 1. Introduction
In recent years, the rapid development of large language models (LLMs) has enabled numerous applications in the legal domain (Brown et al., 2020; Touvron et al., 2023; Scao et al., 2022; Yang et al., 2024), including contract analysis, case law research, and legal document drafting (Shu et al., 2024; Yue et al., 2023; Wang et al., 2024b; Su et al., 2024e). Among these applications, judgment document generation stands out as a particularly promising yet challenging task. Judgment documents are authoritative court records that encapsulate legal reasoning, procedural details, and final rulings (Kumar et al., 2013; Li et al., 2023a). As essential components of judicial proceedings, these documents serve as authoritative records of legal decisions and reflect the complexity of legal reasoning and statutory interpretation (Zander, 2015). In practice, drafting judgment documents is a highly labor-intensive task that consumes substantial time and resources in courts. Specifically, creating a judgment document requires the judge to gather a large amount of legal information, including relevant statutes, past case precedents, and fundamental legal knowledge. Once the relevant legal information has been collected, it must be systematically organized and integrated with professional legal reasoning to generate a well-structured and legally sound judgment document. Due to the time-consuming and labor-intensive nature of drafting judgment documents, AI technologies hold great promise for supporting legal professionals. For example, advanced IR techniques can systematically collect and organize relevant legal information, while generative AI can support drafting well-structured legal texts.
@@ 2.1. Judgment Document Structure and Notation
Let $\mathcal{F}$ denote the space of fact descriptions extracted from judgment documents. Each fact description $f\in\mathcal{F}$ provides a clear and unbiased summary of the circumstances underlying a legal dispute. The corresponding judgment document (denoted by $j$) belongs to the space $\mathcal{J}$. A complete judgment document typically consists of several interdependent sections, as illustrated in Figure 1, including but not limited to:
@@
  • Judicial Reasoning: This section details the court’s evaluation of both factual and legal aspects of the case. It systematically applies relevant laws, legal principles, and judicial precedents to the facts. The analysis discusses the evidence, the offense’s nature, and the liability’s determination. It ensures that the legal reasoning behind the decision is transparent and well-founded.
@@
The interdependence between these sections ensures the judgment document maintains logical consistency. Each section builds upon the previous one, with the Judicial Analysis Process being generated based on the Fact Description while simultaneously providing the legal foundation for the Judgment Result.
Figure 1. An illustration of the structure of a judgment document.
@@ 3. Dataset Construction
This section outlines the construction of the JuDGE dataset, starting with the collection of criminal case documents and the statute corpus. Next, it describes the pre-processing steps to ensure high data quality, followed by expert annotations to verify the dataset’s accuracy and consistency. The section then presents key dataset statistics, offering insights into its composition, and concludes with a discussion of the ethical considerations, highlighting our commitment to privacy, transparency, and regular updates.
@@ 3.3. Expert Annotation
While the data pre-processing steps provided an initial quality filter for judgment documents, we further enhanced the JuDGE dataset through expert annotation. To achieve this, we recruited law students to evaluate each judgment document instance based on five critical aspects: document formatting, the correctness of judicial reasoning, appropriate legal references, the accuracy of crime classification, and the reasonableness of the sentencing. Each document was annotated by two independent experts. If both annotators agreed that the document met the quality criteria, it was classified as “usable.” If discrepancies arose between annotators, the document was excluded from the dataset. The annotation team consisted of seven members recruited from prominent law schools. Our compensation plan, which was based on working hours, ensured that the annotation process was incentivized fairly, offering an average hourly wage of 45 CNY, which is significantly higher than the minimum wage requirements in Beijing. To evaluate the reliability of the annotations, we applied Cohen’s Kappa coefficient in a binary classification context. The analysis, performed on 2,574 annotated instances, yielded a Kappa value of 0.8012, indicating very high inter-annotator agreement. This strong agreement reflects the consistency and reliability of the expert annotations. After applying these filtering standards and expert annotations, we obtained the final dataset with 2,505 legal documents covering 182 different statutes and 142 distinct crime types. This selection process ensured the dataset’s diversity in legal references and crime types while also maintaining its high quality for the benchmark.
@@ 3.4. Dataset Statistics
Table 2 summarizes the core numerical attributes of the JuDGE dataset and its supplementary legal corpora. The dataset contains 2,505 fact-judgment pairs, covering a diverse range of criminal cases with 142 unique charges and 182 distinct criminal law provisions. This diversity suggests that the dataset is highly applicable for modeling various legal scenarios. The average lengths of the different sections further reflect the dataset’s balanced level of detail. Fact descriptions average around 652 Chinese characters, providing a succinct yet informative account of case details. On a per-document basis, the inclusion of roughly 1.26 charges and 4.31 statutory articles emphasizes each case’s multifaceted legal context. Moreover, the dataset is supplemented by two extensive external corpora: one containing 103,251 judgment documents and another comprising 55,348 statutory articles. These resources provide rich legal background information that is particularly valuable for advanced IR technologies like retrieval-augmented generation approaches.
@@ 4. Evaluation Framework
We compare the Judicial Reasoning and Judgment Result sections of the generated judgment document against the ground truth document using the following semantics-based metrics:
@@
While METEOR and BERTScore are not designed specifically for legal reasoning, they offer a practical proxy for evaluating semantic consistency when domain-specific annotations are unavailable. To enhance interpretability and demonstrate the qualitative effectiveness of these metrics, we provide representative case studies on the official dataset website [^4] https://github.com/oneal2000/JuDGE.
@@ 5.1. Baselines
To support future research, we systematically evaluate various baselines on our proposed JuDGE benchmark using different LLMs. These baselines include few-shot in-context learning and fine-tuning techniques to adapt general LLMs to the judgment document generation task. Additionally, we introduce an advanced baseline, Multi-source RAG (MRAG), which integrates knowledge from both statute and judgment document corpora, establishing a strong baseline performance for future reference.
@@ 5.1.3. Retrieval Augmented Generation
Building upon the RAG paradigm, we explore and propose an RAG-based approach as a baseline for our task. In practical legal applications, generating a complete judgment document requires the integration of extensive legal information, including relevant statutory regulations, historical case precedents, and fundamental legal principles. To address this requirement, we propose the Multi-Source Retrieval-Augmented Generation (MRAG) baseline that integrates external knowledge to enhance the generation process. This baseline divides judgment generation into two phases: Information Collection and Document Generation. In the Information Collection phase, two distinct retrievers are employed to retrieve relevant information from different sources. Specifically, the Law Retriever targets relevant statutory regulations, while the Case Retriever focuses on relevant judgment documents. In the Document Generation phase, the LLM leverages the fact description along with the information gathered in the information collection phase to generate the complete judgment document.
@@
Table 4. Experimental results on multiple baselines across six different LLMs. QW-3B-Base refers to the Qwen-2.5-3B model, while QW-3B-Chat refers to the Qwen-2.5-3B-Instruct model (the same naming convention applies to the 7B models). “MET.” denotes METEOR, “BERTS.” for BERTScore, and “Prec.” for precision. The best results for each LLM are highlighted in bold.
@@
Model Method Prison Fine Recall Prec. F1 Recall Prec. F1 MET. BERTS. MET. BERTS.
@@ 6. Experimental Results
(1) Legal LLMs (e.g., LexiLaw-6B and Hanfei-7B) show notably poor performance when generating judgment documents directly or via few-shot in-context learning. A closer examination of the generated outputs reveals that these models often reuse or partially copy the exemplars provided in the prompt, neglecting to incorporate the unique facts of each case. We suspect that this phenomenon arises from catastrophic forgetting. Since these models were continually pre-trained and fine-tuned on large-scale legal data (e.g., legal QA), their parameters have diverged significantly from their initial state, resulting in the loss of previously acquired knowledge and instruction-following capabilities. (2) Few-shot in-context learning benefits larger chat models more significantly. For QW-7B-Chat, introducing a few-shot example in the prompt consistently enhances performance across almost all the metrics. In contrast, QW-3B-Chat sees only marginal gains from in-context learning. This difference highlights that larger models with more capacity generally exhibit stronger in-context learning capabilities. (3) Supervised fine-tuning substantially improves performance. Regardless of the models’ initial capability, supervised fine-tuning (SFT) leads to improvements over direct generation and in-context learning. This trend is observed across both general-purpose models (e.g., QW-3B-Chat and QW-7B-Chat) and legal LLMs (e.g., LexiLaw-6B and Hanfei-7B). Notably, the improvement is particularly large for legal LLMs: while their performance is almost unusable before fine-tuning, it becomes acceptable and sometimes competitive after the task-specific SFT. These findings reinforce the importance of supervised adaptation for certain downstream tasks. (4) Multi-source retrieval-augmented generation (RAG) offers competitive advantages. For instance, MRAG consistently achieves the highest recall, precision, and F1 score in convicting accuracy. However, while MRAG achieves high recall in referencing accuracy, its precision is relatively lower. We attribute this to our top-$k$ retrieval strategy (here $k=10$), which occasionally supplies multiple irrelevant statutes. As indicated by Table 2, only 4.31 criminal statutes are truly relevant per case. Since current LLMs are not fully adept at filtering out misleading articles, this results in a precision drop.
@@ 7.1. Legal Information Retrieval
Legal information retrieval (IR) differs from open-domain IR as it requires the retrieval model to incorporate legal knowledge and consider both semantic and legal perspectives when modeling the text. A prominent research direction is Legal Case Retrieval, which focuses on retrieving past legal cases that are relevant to a given query case, aiding legal professionals in case analysis and decision-making (Su et al., 2023b; Ma et al., 2021, 2023). Early work in this area primarily relied on lexical matching methods to assess similarity (Rosa et al., 2021). Recently, with the development of Dense Retrieval (Karpukhin et al., 2020; Zhan et al., 2021; Su et al., 2023c; Fang et al., 2024; Su et al., 2023a), researchers have begun to explore methods to enhance retrievers with deeper legal understanding via large-scale pretraining on unlabeled legal corpora. For instance, Sailer (Li et al., 2023a) trains the dense retriever through an autoencoder-like approach, where the embedding of a case’s Fact section is used to reconstruct other sections of the document, thereby infusing the fact embeddings with richer contextual information. Caseformer (Su et al., 2023b) proposes an unsupervised contrastive learning approach for automatically measuring similarity between legal documents. Another major direction is Statute Retrieval, which aims to locate and retrieve relevant statutory articles for given a query. For instance, the annual COLIEE competitions (Goebel et al., 2023; Kim et al., 2022; Rabelo et al., 2022; Li et al., 2023b, c) focus on Japanese legal bar exam questions, where the goal is to retrieve relevant statutes from the Japanese Civil Code. Similarly, the AILA (Bhattacharya et al., 2019) targets the Indian legal system, using queries from Supreme Court of India judgments to retrieve relevant statutes. Beyond professional use cases, statute retrieval also caters to non-professional users who may lack formal legal training; for example, STARD (Su et al., 2024a) is specifically designed to address layperson queries, providing plain-language questions paired with relevant statutory content. Beyond retrieval, Legal QA tasks aim to satisfy the public’s need for accessible legal information (Do et al., 2017), such as the French long-form QA dataset LLeQA (Louis et al., 2024) and GerLayQA (Büttner and Habernal, 2024), which contains laymen’s German legal questions paired with authoritative law book paragraphs.
@@ 8. Conclusion
In this paper, we introduced JuDGE, a benchmark for evaluating the generation of judgment documents. By formalizing the task and providing the benchmark dataset, we establish a solid foundation for further research in this important yet underexplored area. Through collaboration with legal professionals, we also developed an automated evaluation framework that measures performance across four key dimensions: penalty accuracy, convicting accuracy, referencing accuracy, and similarity to ground truth. Our experimental results reveal that current approaches struggle to produce high-quality judgment documents. Although supervised fine-tuning and our proposed Multi-source RAG approach show improvements, the performance gap indicates that judgment document generation still remains challenging and requires further research and innovation. We hope JuDGE will inspire more advanced techniques in judgment document generation, ultimately helping the legal community improve efficiency and reduce manual workloads.